Hypothesis Testing

Some people claim that the amount of rainfall this month is very much the same as the great rains of 2010 whose average volume is 450mm with a standard deviation of 100mm. when you checked this month‟s data; you found that the average rainfall from five storms was 180mm. Are the people‟s lay theories correct at α 0.05? Use traditional and p-value method.
1
Expert's answer
2020-05-15T17:59:48-0400

"H_0:\mu=450"

"H_1:\mu\not=450"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test is "z_c=1.96."

The rejection region for this two-tailed test is "R=\{z:|z|>1.96\}"

The z-statistic is computed as follows:


"Z={\bar{X}-\mu\over \sigma/\sqrt{n}}={180-450\over 100/\sqrt{5}}=-6.037"

Using the P-value approach: The p-value is "p=0," and since "p=0<0.05," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 450, at the 0.05 significance level.